You wish to deliver 95 ml of a viscous drug (η = 22×10−4 Pa s) through a needle of length 50 mm and with an internal radius of 0.36 mm.

If a pressure difference of 66 kPa is maintained across the needle as the drug is delivered then how long will it take to deliver the 95 ml of fluid? (to 2 s.f and in s)

(Note: 1 ml = 1×10−6 m3)


1
Expert's answer
2021-04-20T16:49:27-0400

The flow of the fluid can be calculated according to Poiseuille law "Q = \frac{\pi R^4}{8 \eta l}\Delta p".

Hence the time needed to deliver volume "V" of the fluid is:

"t = \frac{V}{Q} = \frac{8 \eta l V}{\pi R^4 \Delta p} \approx 24.01 s".